# refrences.bib

### v1

@@ 2 Preliminary definitions and properties
where $b<u<1$ if $10<b<a<$ 1; $u=b$ if $10<b<a=$ 1; $u=1$ if $a=b>0$; $u<1$ if $a>b=0$; and $u=1$, if $a=b=0$.
It should be noted that if $a>b=0$, we generally have $u\geq 0$. For example, if $\varphi$ is Lukasiewicz t-norm and $a>b=0$, then $u=1-a$ which is positive if $a<1$. However, if $\varphi$ is strict and $a>b=0$, then for each $x>0$, we have $\varphi(a,x)>\varphi(a,0)=0$, i.e., $S=\{0\}$. For each t-norm in Table A1 (see Appendix A), value $u$ discussed in Theorem 1 and Theorem 2 is presented in Table A3.
Definition 2.
@@
According to Definition 2, it is easy to verify that $S_{ij}^{+}\subseteq I_{ij}^{+}$ and $S_{ij}^{-}\subseteq I_{ij}^{-}$, $\forall i\in\mathscr{I}$ and $\forall j\in\mathscr{J}$.
Corollary 2.
@@
    If $a_{ij}^{+}\geq b_{i}=0$, then $S_{ij}^{+}=I_{ij}^{+}=\left[0,u_{ij}^{+}\right]$. Moreover, $u_{ij}^{+}=1$ if $a_{ij}^{+}=b_{i}=0$; and $10\leq u_{ij}^{+}<$ 1 if $a_{ij}^{+}>b_{i}=0$.
@@
    If $a_{ij}^{-}\geq b_{i}=0$, then $S_{ij}^{-}=I_{ij}^{-}=\left[1-u_{ij}^{-},1\right]$. Moreover, $u_{ij}^{-}=1$ if $a_{ij}^{-}=b_{i}=0$; and $10\leq u_{ij}^{-}<$ 1 if $a_{ij}^{-}>b_{i}=0$.
@@
    The results directly follow from Lemma 1 (b), Theorem 1, Theorem 2, Definition 2 and replacing $a_{ij}^{+}$ and $a_{ij}^{-}$ by $a$; $b_{i}$ by $b$; $S_{ij}^{+}$ and $S_{ij}^{-}$ by $S$; and $I_{ij}^{+}$ and $I_{ij}^{-}$ by $I$. ∎
Definition 3.
@@
    (a) The proof is easily resulted from Definitions 2 and 3. (b) According to Definition 3, $x\in S_{ij}$ if and only if $\varphi\left(a_{ij}^{+},x_{j}\right)\leq b_{i}$, $\varphi\left(a_{ij}^{-},1-x_{j}\right)\leq b_{i}$ (i.e., $x\in I_{ij}^{+}\cap I_{ij}^{-}=I_{ij}$) and at least one of the two equalities $\varphi\left(a_{ij}^{+},x_{j}\right)=b_{i}$ (i.e., $\left.x\in S_{ij}^{+}\right)$ and $\left.\varphi\left(a_{ij}^{-},1-x_{j}\right)\right\}=b_{i}$ (i.e., $x\in S_{ij}^{-}$) holds. ∎
Based on Corollary 2 and Lemma 3, we can obtain the following results, which characterize sets $S_{ij}$ and $I_{ij}$ for all cases.
Corollary 3.
@@
$S_{ij}=\left\{\begin{array}[]{ll}\varnothing&,a_{ij}^{+},a_{ij}^{-}<b_{i}\\ \left\{u_{ij}^{+}\right\}&,a_{ij}^{+}\geq b_{i}>0\text{ and }a_{ij}^{-}<b_{i}\\ \left\{1-u_{ij}^{-}\right\}&,a_{ij}^{-}\geq b_{i}>0\text{ and }a_{ij}^{+}<b_{i}\\ {\left[1-u_{ij}^{-},u_{ij}^{+}\right]}&,a_{ij}^{+},a_{ij}^{-}\geq b_{i}=0\\ \left\{1-u_{ij}^{-},u_{ij}^{+}\right\}&,a_{ij}^{+},a_{ij}^{-}\geq b_{i}>0\end{array}\quad,\quad I_{ij}=\begin{cases}{[0,1]}&,a_{ij}^{+},a_{ij}^{-}<b_{i}\\ {\left[0,u_{ij}^{+}\right]}&,a_{ij}^{+}\geq b_{i}>0\text{ and }a_{ij}^{-}<b_{i}\\ {\left[1-u_{ij}^{-},1\right]}&,a_{ij}^{-}\geq b_{i}>0\text{ and }a_{ij}^{+}<b_{i}\\ {\left[1-u_{ij}^{-},u_{ij}^{+}\right]}&,a_{ij}^{+},a_{ij}^{-}\geq b_{i}\geq 0\end{cases}\right.$
where $b_{i}<u_{ij}^{+}<1$ if $10<b_{i}<a_{ij}^{+}<$ 1; $u_{ij}^{+}=b_{i}$ if $10<b_{i}<a_{ij}^{+}=$ 1; $u_{ij}^{+}=1$ if $a_{ij}^{+}=b_{i}$ ; $10\leq u_{ij}^{+}<$ 1 if $a_{ij}^{+}>b_{i}=0$; $b_{i}<u_{ij}^{-}<1$ if $10<b_{i}<a_{ij}^{-}<$ 1; $u_{ij}^{-}=b_{i}$ if $10<b_{i}<a_{ij}^{-}=$ 1; $u_{ij}^{-}=1$ if $a_{ij}^{-}=b_{i}$; and $10\leq u_{ij}^{-}<$ 1 if $a_{ij}^{-}>b_{i}=0$.
Remark 3.
According to Remark 2, each $I_{ij}$ can be written as an interval $I_{ij}=\left[L_{ij},U_{ij}\right]$ where $L_{ij}$ and $U_{ij}$ are the lower and upper bounds of the interval, respectively. Also, each $S_{ij}$ (if it is non-empty) is obtained as one of the forms of $S_{ij}=\left\{L_{ij}\right\}$, $S_{ij}=\left\{U_{ij}\right\}$, $S_{ij}=\left\{L_{ij},U_{ij}\right\}$ or $S_{ij}=\left[L_{ij},U_{ij}\right]$.
Definition 4.
@@
    From Definition 4 and Remark 3, we have
@@
    Hence, for each $i\in\mathscr{I}$, if $L_{ij}\in I_{j}$ $\left(U_{ij}\in I_{j}\right)$, then we must have $L_{ij}=L_{j}$ $\left(U_{ij}=U_{j}\right)$. On the other hand, from Remark 3, a non-empty $S_{ij}$ is as one of the forms of $S_{ij}=\left\{L_{ij}\right\}$, $S_{ij}=\left\{U_{ij}\right\}$, $S_{ij}=\left\{L_{ij},U_{ij}\right\}$ or $S_{ij}=\left[L_{ij},U_{ij}\right]$. So, if $S_{ij}=\left\{L_{ij}\right\}$ and $S_{ij}^{\prime}=S_{ij}\cap I_{j}\neq\varnothing$, then $L_{ij}\in I_{j}$; that is, $L_{ij}=L_{j}$ and therefore $S_{ij}^{\prime}=S_{ij}\cap I_{j}=\left\{L_{j}\right\}$. For the other cases, the results are attained by the similar argument. ∎
Example 1.
@@
In this example, =$\mathscr{I}=\{1,2,\ldots,10\}$, $\mathscr{J}=\{1,2,\ldots,10\}$ and $\varphi$ is the Yager t-norm (that is continuous and nilpotent) with $p=2$. So, according to Table A1 we have
@@
Also, $f_{Y}(x)=(1-x)^{2}$ and $f_{Y}^{(-1)}(x)=\max\{1-\sqrt{x},0\}$ (see Table A2). For $i=8$ and $j=9$, we have $a_{89}^{+}=1$ and $a_{89}^{-}=b_{8}=0.8$. Thus, from Corollary 3 (Part (e)) we obtain $S_{89}=\left\{1-u_{89}^{-},u_{89}^{+}\right\}$ and $I_{89}=\left[1-u_{89}^{-},u_{89}^{+}\right]$, where $u_{89}^{-}$ and $u_{89}^{+}$ are calculated by $u_{89}^{-}=1-\sqrt{\left(1-b_{8}\right)^{2}-\left(1-a_{89}^{-}\right)^{2}}=1$ and $u_{89}^{+}=1-\sqrt{\left(1-b_{8}\right)^{2}-\left(1-a_{89}^{+}\right)^{2}}=0.8$ (see Table A3). Therefore, $S_{89}=\{0,0.8\}$ and $I_{89}=[0,0.8]$. Also, see Remark 2, that indicates $u_{ij}^{+}=b_{i}$ if $10<b_{i}<a_{ij}^{+}=$ 1 and $u_{ij}^{-}=1$ if $a_{ij}^{-}=b_{i}$ for any arbitrary continuous Archimedean t-norm. By the similar calculations, it follows that $I_{i9}=[0,1]$ for $i\in\{1,\ldots,5\}\bigcup\{7,9,10\}$ and $I_{69}=[0.2,0.8]$. So, from Definition 4, we have $I_{9}=\bigcap_{i\in\mathscr{I}}I_{i9}=[0.2,0.8]$ and $S_{89}^{\prime}=S_{89}\cap I_{9}=\{0.8\}$. In Section 6, Tables $1-4$ show all sets $I_{ij}$, $S_{ij},I_{j}$ and $S_{ij}^{\prime}$ for each $i\in\mathscr{I}$ and $j\in\mathscr{J}$, respectively.
@@ 3 Resolution of the feasible solution set for Problem (1)
    (a) By contradiction, suppose that $S\left(A^{+},A^{-},b\right)\neq\varnothing$ and $I_{j_{0}}=\varnothing$ for some $j_{0}\in\mathscr{J}$. Hence, there exists $i_{0}\in\mathscr{I}$ such that $x\notin I_{i_{0}j_{0}}$ (Definition 4), $\forall x\in S$ $\left(A^{+},A^{-},b\right)$. So, from Definition 3, we have
@@
    that implies $x\notin S_{i_{0}}$. Consequently, since $S\left(A^{+},A^{-},b\right)=\bigcap_{i\in\mathscr{I}}S_{i}$, we obtain $x\notin S\left(A^{+},A^{-},b\right)$ that is a contradiction. (b) By contradiction, assume that $x\in S$ $\left(A^{+},A^{-},b\right)$ and there exists some $i_{0}\in\mathscr{I}$ such that $S_{i_{0}j}^{\prime}=\varnothing$, $\forall j\in\mathscr{J}$. So, $x_{j}\notin S_{i_{0}j}^{\prime}=S_{i_{0}j}\cap I_{j}$, $\forall j\in\mathscr{J}$ (Definition 4). Now, if $x_{j}\notin I_{j}$ for some $j\in\mathscr{J}$, then from Part (a) we obtain $S\left(A^{+},A^{-},b\right)\neq\varnothing$, that is a contradiction. On the other hand, if $x_{j}\notin S_{i_{0}j}^{\prime}$ and $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$, then it is concluded that $x_{j}\notin S_{i_{0}j}$, $\forall j\in\mathscr{J}$, that implies $\max\left\{\varphi\left(a_{i_{0}j}^{+},x_{j}\right),\varphi\left(a_{i_{0}j}^{-},1-x_{j}\right)\right\}<b_{i_{0}}$, $\forall j\in\mathscr{J}$ (Definition 3). Therefore, $x\notin S_{i_{0}}$, that contradicts the assumption that $x\in S$ $\left(A^{+},A^{-},b\right)$. ∎
@@
Remark 4.
From Lemma 5, it follows that if $x\in S\left(A^{+},A^{-},b\right)$, then $L_{j}\leq x_{j}\leq U_{j}$, $\forall j\in\mathscr{J}$ , where $L_{j}$ and $U_{j}$ are the lower and upper bound of $I_{j}=\left[L_{j},U_{j}\right]$, respectively.
Corollary 5.
@@
    The proof is resulted from Lemma 5, where $S\left(A^{+},A^{-},b\right)$ and $I_{j}$ are replaced by $S_{i}$ and $I_{ij}$, respectively. ∎
Definition 5.
@@
Remark 5.
According to Corollary 6, the number of admissible functions is bounded above by $\prod_{i\in\mathscr{I}}\left|\mathscr{J}_{i}\right|$, where $\left|\mathscr{J}_{i}\right|$ denotes the cardinality of $\mathscr{J}_{i}$. However, the actual number of admissible functions is usually much less than this value.
@@
According to (11), for each $e\in E,~{}x\in S(e)$ and $j\in\mathscr{J}$, we have either $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}$ (if $\mathscr{I}_{j}(e)\neq\varnothing$ ) or $x_{j}\in I_{j}=\left[L_{j},U_{j}\right]$ (if $\mathscr{I}_{j}(e)=\varnothing$ ). On the other hand, in the former case, we have $\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}\neq\varnothing$ (from Corollary 6) and also $\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}$ is as one of the forms of $\left\{L_{j}\right\},~{}\left\{U_{j}\right\},~{}\left\{L_{j},U_{j}\right\}$ or $\left[L_{j},U_{j}\right]$ (from Corollary 4). So, it is concluded that $S(e)$ ( $\forall e\in E$ ) is indeed a closed-bounded subset (and therefore, a compact subset) of $[0,1]^{n}$. This fact together with Theorem 3 implies that the feasible solution set of Problem (1) is determined by the union of a finite number of compact sets.
Example 2.

### v2

@@ 2 Preliminary definitions and properties
where $b<u<1$ if $0<b<a<1$; $u=b$ if $0<b<a=1$; $u=1$ if $a=b>0$; $u<1$ if $a>b=0$; and $u=1$, if $a=b=0$.
It should be noted that if $a>b=0$, we generally have $u\geq 0$. For example, if $\varphi$ is Lukasiewicz t-norm and $a>b=0$, then u=1au=1-a which is positive if $a<1$. However, if $\varphi$ is strict and $a>b=0$, then for each $x>0$, we have $\varphi(a,x)>\varphi(a,0)=0$, i.e., $S=\{0\}$. For each t-norm in Table A1 (see Appendix A), value $u$ discussed in Theorem 2.5 and Theorem 2.6 is presented in Table A3.
Definition 2.8.
@@
According to Definition 2.8, it is easy to verify that $S_{ij}^{+}\subseteq I_{ij}^{+}$ and $S_{ij}^{-}\subseteq I_{ij}^{-}$, $\forall i\in\mathscr{I}$ and $\forall j\in\mathscr{J}$.
Corollary 2.9.
@@
    If $a_{ij}^{+}\geq b_{i}=0$, then $S_{ij}^{+}=I_{ij}^{+}=\left[0,u_{ij}^{+}\right]$. Moreover, $u_{ij}^{+}=1$ if $a_{ij}^{+}=b_{i}=0$; and $0\leq u_{ij}^{+}<1$ if $a_{ij}^{+}>b_{i}=0$.
@@
    If $a_{ij}^{-}\geq b_{i}=0$, then $S_{ij}^{-}=I_{ij}^{-}=\left[1-u_{ij}^{-},1\right]$. Moreover, $u_{ij}^{-}=1$ if $a_{ij}^{-}=b_{i}=0$; and $0\leq u_{ij}^{-}<1$ if $a_{ij}^{-}>b_{i}=0$.
@@
The results directly follow from Lemma 2.2 (b), Theorem 2.5, Theorem 2.6, Definition 2.8 and replacing $a_{ij}^{+}$ and $a_{ij}^{-}$ by $a$; $b_{i}$ by $b$; $S_{ij}^{+}$ and $S_{ij}^{-}$ by $S$; and $I_{ij}^{+}$ and $I_{ij}^{-}$ by $I$. ∎
Definition 2.10.
@@
(a) The proof is easily resulted from Definitions 2 and 3. (b) According to Definition 2.10, $x\in S_{ij}$ if and only if $\varphi\left(a_{ij}^{+},x_{j}\right)\leq b_{i}$, $\varphi\left(a_{ij}^{-},1-x_{j}\right)\leq b_{i}$ (i.e., $x\in I_{ij}^{+}\cap I_{ij}^{-}=I_{ij}$) and at least one of the two equalities $\varphi\left(a_{ij}^{+},x_{j}\right)=b_{i}$ (i.e., $\left.x\in S_{ij}^{+}\right)$ and $\left.\varphi\left(a_{ij}^{-},1-x_{j}\right)\right\}=b_{i}$ (i.e., $x\in S_{ij}^{-}$) holds. ∎
Based on Corollary 2.9 and Lemma 2.11, we can obtain the following results, which characterize sets $S_{ij}$ and $I_{ij}$ for all cases.
Corollary 2.12.
@@
[…]
where $b_{i}<u_{ij}^{+}<1$ if $0<b_{i}<a_{ij}^{+}<1$; $u_{ij}^{+}=b_{i}$ if $0<b_{i}<a_{ij}^{+}=1$; $u_{ij}^{+}=1$ if $a_{ij}^{+}=b_{i}$ ; $0\leq u_{ij}^{+}<1$ if $a_{ij}^{+}>b_{i}=0$; $b_{i}<u_{ij}^{-}<1$ if $0<b_{i}<a_{ij}^{-}<1$; $u_{ij}^{-}=b_{i}$ if $0<b_{i}<a_{ij}^{-}=1$; $u_{ij}^{-}=1$ if $a_{ij}^{-}=b_{i}$; and $0\leq u_{ij}^{-}<1$ if $a_{ij}^{-}>b_{i}=0$.
Remark 2.14.
According to Remark 2.13, each $I_{ij}$ can be written as an interval $I_{ij}=\left[L_{ij},U_{ij}\right]$ where $L_{ij}$ and $U_{ij}$ are the lower and upper bounds of the interval, respectively. Also, each $S_{ij}$ (if it is non-empty) is obtained as one of the forms of $S_{ij}=\left\{L_{ij}\right\}$, $S_{ij}=\left\{U_{ij}\right\}$, $S_{ij}=\left\{L_{ij},U_{ij}\right\}$ or $S_{ij}=\left[L_{ij},U_{ij}\right]$.
Definition 2.15.
@@
From Definition 2.15 and Remark 2.14, we have
@@
Hence, for each $i\in\mathscr{I}$, if $L_{ij}\in I_{j}$ $\left(U_{ij}\in I_{j}\right)$, then we must have $L_{ij}=L_{j}$ $\left(U_{ij}=U_{j}\right)$. On the other hand, from Remark 2.14, a non-empty $S_{ij}$ is as one of the forms of $S_{ij}=\left\{L_{ij}\right\}$, $S_{ij}=\left\{U_{ij}\right\}$, $S_{ij}=\left\{L_{ij},U_{ij}\right\}$ or $S_{ij}=\left[L_{ij},U_{ij}\right]$. So, if $S_{ij}=\left\{L_{ij}\right\}$ and $S_{ij}^{\prime}=S_{ij}\cap I_{j}\neq\varnothing$, then $L_{ij}\in I_{j}$; that is, $L_{ij}=L_{j}$ and therefore $S_{ij}^{\prime}=S_{ij}\cap I_{j}=\left\{L_{j}\right\}$. For the other cases, the results are attained by the similar argument. ∎
Example 2.17.
@@
In this example, $\mathscr{I}=\{1,2,\ldots,10\}$, $\mathscr{J}=\{1,2,\ldots,10\}$ and $\varphi$ is the Yager t-norm (that is continuous and nilpotent) with $p=2$. So, according to Table A1 we have
@@
Also, $f_{Y}(x)=(1-x)^{2}$ and $f_{Y}^{(-1)}(x)=\max\{1-\sqrt{x},0\}$ (see Table A2). For $i=8$ and $j=9$, we have $a_{89}^{+}=1$ and $a_{89}^{-}=b_{8}=0.8$. Thus, from Corollary 2.12 (Part (e)) we obtain $S_{89}=\left\{1-u_{89}^{-},u_{89}^{+}\right\}$ and $I_{89}=\left[1-u_{89}^{-},u_{89}^{+}\right]$, where $u_{89}^{-}$ and $u_{89}^{+}$ are calculated by $u_{89}^{-}=1-\sqrt{\left(1-b_{8}\right)^{2}-\left(1-a_{89}^{-}\right)^{2}}=1$ and $u_{89}^{+}=1-\sqrt{\left(1-b_{8}\right)^{2}-\left(1-a_{89}^{+}\right)^{2}}=0.8$ (see Table A3). Therefore, $S_{89}=\{0,0.8\}$ and $I_{89}=[0,0.8]$. Also, see Remark 2.13, that indicates $u_{ij}^{+}=b_{i}$ if $0<b_{i}<a_{ij}^{+}=1$ and $u_{ij}^{-}=1$ if $a_{ij}^{-}=b_{i}$ for any arbitrary continuous Archimedean t-norm. By the similar calculations, it follows that $I_{i9}=[0,1]$ for $i\in\{1,\ldots,5\}\bigcup\{7,9,10\}$ and $I_{69}=[0.2,0.8]$. So, from Definition 2.15, we have $I_{9}=\bigcap_{i\in\mathscr{I}}I_{i9}=[0.2,0.8]$ and $S_{89}^{\prime}=S_{89}\cap I_{9}=\{0.8\}$. In Section 6, Tables 141-4 show all sets $I_{ij}$, $S_{ij},I_{j}$ and $S_{ij}^{\prime}$ for each $i\in\mathscr{I}$ and $j\in\mathscr{J}$, respectively.
@@ 3 Resolution of the feasible solution set for Problem (1)
(a) By contradiction, suppose that $S\left(A^{+},A^{-},b\right)\neq\varnothing$ and $I_{j_{0}}=\varnothing$ for some $j_{0}\in\mathscr{J}$. Hence, there exists $i_{0}\in\mathscr{I}$ such that $x\notin I_{i_{0}j_{0}}$ (Definition 2.15), $\forall x\in S$ $\left(A^{+},A^{-},b\right)$. So, from Definition 2.10, we have
@@
that implies $x\notin S_{i_{0}}$. Consequently, since $S\left(A^{+},A^{-},b\right)=\bigcap_{i\in\mathscr{I}}S_{i}$, we obtain $x\notin S\left(A^{+},A^{-},b\right)$ that is a contradiction. (b) By contradiction, assume that $x\in S$ $\left(A^{+},A^{-},b\right)$ and there exists some $i_{0}\in\mathscr{I}$ such that $S_{i_{0}j}^{\prime}=\varnothing$, $\forall j\in\mathscr{J}$. So, $x_{j}\notin S_{i_{0}j}^{\prime}=S_{i_{0}j}\cap I_{j}$, $\forall j\in\mathscr{J}$ (Definition 2.15). Now, if $x_{j}\notin I_{j}$ for some $j\in\mathscr{J}$, then from Part (a) we obtain $S\left(A^{+},A^{-},b\right)\neq\varnothing$, that is a contradiction. On the other hand, if $x_{j}\notin S_{i_{0}j}^{\prime}$ and $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$, then it is concluded that $x_{j}\notin S_{i_{0}j}$, $\forall j\in\mathscr{J}$, that implies $\max\left\{\varphi\left(a_{i_{0}j}^{+},x_{j}\right),\varphi\left(a_{i_{0}j}^{-},1-x_{j}\right)\right\}<b_{i_{0}}$, $\forall j\in\mathscr{J}$ (Definition 2.10). Therefore, $x\notin S_{i_{0}}$, that contradicts the assumption that $x\in S$ $\left(A^{+},A^{-},b\right)$. ∎
@@
Remark 3.3.
From Lemma 3.2, it follows that if $x\in S\left(A^{+},A^{-},b\right)$, then $L_{j}\leq x_{j}\leq U_{j}$, $\forall j\in\mathscr{J}$ , where $L_{j}$ and $U_{j}$ are the lower and upper bound of $I_{j}=\left[L_{j},U_{j}\right]$, respectively.
Corollary 3.4.
@@
The proof is resulted from Lemma 3.2, where $S\left(A^{+},A^{-},b\right)$ and $I_{j}$ are replaced by $S_{i}$ and $I_{ij}$, respectively. ∎
Definition 3.5.
@@
Remark 3.9.
According to Corollary 3.8, the number of admissible functions is bounded above by $\prod_{i\in\mathscr{I}}\left|\mathscr{J}_{i}\right|$, where $\left|\mathscr{J}_{i}\right|$ denotes the cardinality of $\mathscr{J}_{i}$. However, the actual number of admissible functions is usually much less than this value.
@@
According to (11), for each $e\in E,~x\in S(e)$ and $j\in\mathscr{J}$, we have either $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}$ (if $\mathscr{I}_{j}(e)\neq\varnothing$ ) or $x_{j}\in I_{j}=\left[L_{j},U_{j}\right]$ (if $\mathscr{I}_{j}(e)=\varnothing$ ). On the other hand, in the former case, we have $\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}\neq\varnothing$ (from Corollary 3.8) and also $\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}$ is as one of the forms of $\left\{L_{j}\right\},~\left\{U_{j}\right\},~\left\{L_{j},U_{j}\right\}$ or $\left[L_{j},U_{j}\right]$ (from Corollary 2.16). So, it is concluded that $S(e)$ ( $\forall e\in E$ ) is indeed a closed-bounded subset (and therefore, a compact subset) of $0,1]^{n}$. This fact together with Theorem 3.10 implies that the feasible solution set of Problem (1) is determined by the union of a finite number of compact sets.
Example 3.11.